A has 3 shares in a lottery contain 3 prizes and 9 blanks & B has 2 shares in a lottery contain 2 prizes & 6 blanks Compare chances of winning ?


 

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mayank

probability for A to loss = 9C3 /12C3

probability for A to win  = 1 - 9C3 /12C3  =1 -84/220  =136/220

 

probability for B to loss = 6C2 /8C2

probability for A to win  = 1 - 6C2 /8C2  =1 -15/28  =13/28

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin